OpenStudy (anonymous):

consider he right triangle ABC given below: Part1: use the degree angle and the sine ratio to find the length of side b to two decimal places. Part2: you can find the length of side a using cos25degree, tan25degree or the pythagorean theorem. fid the length of side a 2 decimal places sing one of these methods

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sin 25 = \frac{b}{25} \implies 25 * \sin 25 = b\]

OpenStudy (anonymous):

is it -3.330

OpenStudy (anonymous):

what's sin 25?

OpenStudy (anonymous):

-.323517501

OpenStudy (anonymous):

is angle ABC 25 degrees?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it can't be 25 since \[\cos 25 \neq 9/25\]

OpenStudy (anonymous):

what did you end up getting for the answers @tdiaz100